The company is planning to acquire new equipment for a cost of $38000. Management has already conducted and paid $15000 for a marketing survey while delivery and installation costs are expected to be $9000 and $7400 respectively. Financing costs are $3000 per year. The economic life of the investment in 5 years and the new equipment will depreciate straight-line to a zero value over 5 years. The management learn expects this new equipment at the end of 5 years at a price of $4000

It expects the acquisition to increase sales by $6000 annually while cash operation costs are expected to increase by $3200 annually . This investment will allow the replacement of old equipment which can be sold for $5000. It had a book value of $10000 four years ago, but it has now fully depreciated. The marginal tax rate is 25% and the expected rate of return is 12%

Additional net working capital of $8000 will be needed immediately. When the project is terminated in 5 years, there no longer will be a need for this incremental working capital

(a)Compute the net investment

(b)Compute the net cash flow for each of the five years of the project (year 1-5)

(c)What is the NPV of the project
Answer #1

a) Statement showing net investment

ParticularsAmount
Cost of new equipment38000
Add: Delivery cost9000
Add: Installation cost7400
Add: Wc required8000
Less: Cash flow from the sale of old equipment
= 5000 (1-tax rate)
= 5000(1-25%)
= 5000(0.75)
= 3750
3750
Net investment58650

Thus net investment = $58650

b) Statement showing OFC

Year12345
Increase in sales60006000600060006000
Less: Expense32003200320032003200
Less: Depreciation1088010880108801088010880
PBT-8080-8080-8080-8080-8080
Tax @ 25%-2020-2020-2020-2020-2020
PAT-6060-6060-6060-6060-6060
Add: Depreciation1088010880108801088010880
OCF48204820482048204820

Statement showing total cash flow

Year12345Total
OCF48204820482048204820
Terminal cash flow



11000
Total cash flow482048204820482015820

Note:

Cash flow from the sale of old equipment

ParticularsAmount
The selling price of equipment at end of year 54000
Book value of equipment0
Profit4000
Tax @ 25%1000
Cash flow from the sale of equipment
(4000-1000)
3000

Statement showing Terminal cash flow

ParticularsAmount
Cash flow from the sale of machine3000
Release of WC8000
Terminal cash flow11000

c) Statement showing NPV

Year12345Total
OCF48204820482048204820
Terminal cash flow



11000
Total cash flow482048204820482015820
PVIF @ 12%0.89290.79720.71180.63550.5674
PV4303.573842.473430.783063.208976.6923616.72
Less: Initial Investment




58650.00
NPV




-35033.28

Thus NPV = -35033.28 $


answered by: qbank